Forward Silicon Tracker Module Production and Readout Test at Fermi lab

Test Stand Setup


01/05/2021: Setup MPOD Crate
Installed the ARC (took from STAR FSTMPOD01 RDO2) on the 7th slot of WIENER MPOD Crate
Installed ARM (#49 took from STAR FSTMPOD01 RDO1) on the 2nd slot of WIENER MPOD Crate
Installed ABC (from Mike) on the 2nd slot of WIENER MPOD Crate.
Connected the HV cable to the ABC board: u0 connected to the bottom BNC connector (outer sensor) on the ABC board and u1 connected to the bottom BNC-pin connector (inner sensor) on the ABC board.
Connected gray cable to the bottom slot on ABC board.
Connected the power cable for the WIENER MPOD Crate.

    


01/14/2021: Setup DarkBox and Cable
Connect FST-V3.1-01 to FST04 (prototype).
Place FST04 with purple cable connected inside the dark box.
Mount purple cable into dark box (left photo).
Connect purple cable and gray cable to the PPB (middle photo).
Test stand setup at 01/14/2021 can be viewed from the right photo.

    

01/14/2021: DAQ Driver Installation
Modify /boot/grub2/grub.cfg and reboot the fst-uic machine.
resmem.ko & rorc.ko driver installation failed due to the difference in the kernel version.


01/15/2021: DAQ Driver Installation
Successfully installed new resmem.ko driver from Tonko.


01/16/2021: DAQ Driver Installation
Successfully installed new rorc_driver.ko.
Successfully run how_to.sh.
DAQ installation complete & need to verify with readout test.
The DAQ execs are under: /home/xusun/FstFermiLabTestStand
- Uploaded new fgt_run_daq & fgt_read_daq (from Tonko) to fst-uic.
To run the daq:
- ./fgt_run_daq -X fst -R e1 -L -c FST_conf_UIC_Outer.txt -n 100 -S ./Data/test.sfs
- Use -X fst to indicate the output with fst format → without -X option is to use old fgt format
- Configuration files put under: /RTS/conf/fst/
  - FST_conf_UIC_Outer.txt
  - FST_conf_BNL.txt


01/18/2021: Finalize FNAL Test Stand
Mount PC on the Crate.
Setup Pulse Generator on top of the PC and connect the power cable of Pulse Generator to the rack.
Connect ARC and DAQ PC with optical fiber.
Re-organize the cables.
Clean up the desk.
    


01/19/2020: Setup Slow Control System
Installed SNMP protocol with the help of Scott Reid.


01/20/2020: Finize FNAL Test Stand
Remove two unused ARM modules from the MPOD crate.
Connect Pulse Generator to the ARC as event trigger.
Further clean the desk setup.
Finalize the documents for the FNAL inspection.
  


01/27/2021: ORC Review
ORC Review happened at 3 pm.
No big issue found and additional support documents are required.


01/29/2021: Clean up the Test Stand
Reconnect & organised cable based on the recomandation from ORC review.
Couldn't log into fst-uic machine after reboot.
Contact Scott for help.


01/31/2021: ORC Approved
Got approved from ORC at 20:15.


02/06/2021: Access MPOD from mac
The static IP address of MPOD controller is 192.168.75.101.
To access the MPODc from mac
- Change IPV4 address to 192.168.75.100
- Change subnet mask to 255.255.0.0
- Change Router to 192.168.75.0
- Then LAN connection should turn green.
Open a browser and tap MPODc IP address, the following page should appear and refreshing every few second:



02/07/2021: Able to read out FST04 at SiDet
Got new fgt_run_daq from Tonko → could readout temperature sensor on FST module.
The pulse generator is set to a square waveform with 1V amplitude and -500mV offset at 1KHz frequency.
- need to check the waveform under a scope to make sure the amplitude to be sent to ARC.
ARC can accept the trigger and readout FST04 with VPSP = 38.
FST04 behaves the same as the test in UIC/BNL.

FST Test Stand at SiDet is working now.


02/09/2021: Able to control MPODc remotely
Put WIENER-CRATE-MIB.txt to /usr/share/snmp/mibs.
Set the IP address as described in 02/06/2021.
All snmp command should work.


02/10/2021: Implemented Slow Control Script on Mac
Tested FstOn.sh FstOff.sh FstSD.sh → all works as expected.
Adjust the current limit from 1000 muA to 10 muA → typo from previous IST SC.
The code is under.


02/11/2021: Finalize Slow Control Script
FstOn.sh: Turn ON the crate and configure.
FstRampHV.sh: Ramp to any HV.
FstOff.sh: Ramp down the HV (if not 0) and turn OFF the crate.


02/13/2021: Test Stand Investigation
Change ABC12 to ABC26 board: u0 connected to the bottom BNC connector (outer sensor) on the ABC board and u1 connected to the bottom BNC-pin connector (inner sensor) on the ABC board.
Change GC26 to GC13: fixed the noise issue on APV5.
Put ABC12 back and no more noise issue on APV5.



FST Module Assembly
Detailed information for each module can be found at: http://leto.phy.uic.edu:8080/FST/
Below is only the general log for module assembly at FNAL. 


01/25/2021: FST-28 & FST-33 Assembly Start
Put FST-28 in Storage Box #1 & FST-33 in Storage Box #2.
Put the FST-28 & FST-33 in the transfer area in SiDet.


01/27/2021: FST-28 & FST-33 RO Test at UIC
Perform RO test at UIC.
FST-28 behaves as expected.
FST-33 has no readout when both inner & outer are connected.
- Could read out inner sector.
- Expect a short in the outer sector.
- Found some wires in the APV Chip 6.


01/28/2021: FST-33 RO Test at UIC
Perform RO test at UIC again after Fermilab Technician removed the wires.
Results stay the same as last test.

01/28/2021: FST-21 & FST-24 Assembly Start
Perform Open/Short test on FST-21 & FST-24.
- NO issue identified.
Put FST-21 in Storage Box #3 and FST-24 in Storage Box #4.
Put FST-21 & FST-24 in transfer area in SiDet.


01/29/2021: FST-33 Check
Identified short between VSSa (pin 1 & 3 & 5) and GND (pin 7 & 9 & 11).
Inspect APV chips wire bonding at Willson Hall.
- NO open/short connection found.
Spot a possible soldering wire under pin 5 & 7.
Made open/short test a standard operation procedure for future assembly.


02/01/2021: Investigation of Short between VSSa & GND
Received FST-21 & FST-24 at SiDet. Both of them show a short between FST-21 & FST-24.
Turns out to be silver glue overflowed from VSSa to the GND for both FST-21 & FST-24. Therefore, a short happened.
FST-28 & FST-33 have the same issue. Luckily FST-28 didn't have the short.
Bert fixed the epoxy issue for all 4 modules.
No short between VSSa & GND anymore.


02/01/2021 Open & Short Test Investigation
Test FST-12 & FST-18 & FST-31.
No open and short connection found.

02/01/2021: FST-18 & FST-31 Assembly Start
Put FST-18 in Storage Box #5 and FST-31 in Storage Box #6.
Put FST-18 & FST-31 in transfer area in SiDet.


02/02/2021 Open & Short Test Investigation
Test FST-09 & FST-11 & FST-29 
& FST-30 & FST-34 & FST-35.
No open and short connection found.

02/02/2021: RO Test for FST-28 & FST-33 & FST-21 & FST-24
Test performed at UIC.
The short between VSSa and GND is fixed for all four modules and method is working.
One chip missing (APV5) on FST-21. Normal Pedestal and Noise for the rest of modules.
Put FST-33 to the transfer area to fix the broken wires caused by the attempt to fix the epoxy issue.
An open is observed between pin OUT0_P (14) and OUT0_N (16) => under investigation.


02/03/2021: FST-34 & FST-35 & FST-30 & FST-09 Assembly Start
Put FST-34 in Storage Box #7, FST-35 in Storage Box #8, FST-30 in Storage Box #9 and FST-09 in Storage Box #10. FST-30 and FST-09 are slightly suppressed by the top cover of storage box.

02/03/2021: RO Test for FST-18 & FST-31
Inspect FST18 & FST-31 and repaired FST-33. No issue found.
RO Test performed at UIC.
Normal Pedestal and Noise for all of the modules. Need to change VPSP to 38 for FST-18 and FST-31.


02/05/2021: RO Test for FST-34 & FST-35
Inspect FST-34 and FST-35. The cooper layer of GND on outer sector was removed by Fermilab Technician.
The Cut on FST-35 & APV 6 is deeper than others and some remaining material is around.
Test performed at UIC.
Normal pedestal and noise for all of the modules. Need to change VPSP to 38 for FST-35 to fix the low pedestal and noise channel. Both VPSP41 and 38 works for FST-34.


02/06/2021 Open & Short Test Investigation
Test FST-37 & FST-38
.
No open and short connection found.

02/06/2021: FST-38 Assembly Start
Put FST-38 in Storage Box #11. 


02/08/2021: FST-37 Assembly Start
Put FST-37 in Storage Box #12. 
Put FST-37 & FST-38 in transfer area in SiDet.

02/08/2021 Open & Short Test Investigation
Test FST-45 
& FST-22 & FST-47 & FST-43 & FST-42 & FST-36 & FST-41 & FST-50.
No open and short connection found.

02/08/2021: RO Test for FST-09 & FST-30
Inspect FST-09 and FST-30. The cooper layer of GND on outer sector was removed by Fermilab Technician.
No short found. 
Test performed at FermiLab.
Normal pedestal and noise for FST-30.
No readout of pedestal and noise for both inner and outer sectors of FST-09 => under investigation.
No readout of temperature sensors for both inner and outer sectors of FST-09 => under investigation.

02/09/2021: APV RO Test for FST-38 & FST-37
Inspect FST-38 and FST-37. The cooper layer of GND on outer sector was removed by Fermilab Technician.  
No short found. 
The wire 13 and 14 are short on APV 5 for FST 37.
Test performed at FermiLab.
Normal pedestal and noise for FST-38.
APV 5 is missing for FST-37, which may be caused by the short connection between wire 13 and 14. Need to double check and remake the RO Test after repairing the issue.

02/10/2021: Si Sensor RO test for FST-18
Inner sensor mounted.
Normal pedestal and noise with HV.
Issue: The noise of APV (w/o si sensor) on outer sector got affected with HV applied → under investigation.


02/11/2021: Si Sensor RO test for FST-31
Outer sensor mounted.
Normal pedestal and noise with HV.
APV on inner sector behaves normal with different HV applied.
Start to apply HV to all module with APV only to check if any issue identified.
FST28: No Issue.
FST33: No Issue.
FST24: No Issue.
FST34: No Issue.
F
ST35: No Issue.
FST30: No Issue.
FST38: No Issue.


02/12/2021: FST-18 Inner Sensor RO test at UIC
Normal pedestal and noise with different HV for inner sensor of FST-18.
No noise issue on APV5 at different HV.
Suspect SiDet test stand has some noise issue → bring all the relevent part from UIC to SiDet for cross check.


02/13/2021: FST-18 Inner Sensor RO test at SiDet
GC26 is identified to be problematic.
After switch GC26 to GC13, NO noise issue anymore on APV5.


02/17/2021:FST-18 Outer Sensor and FST-31 Inner Sensor RO test at SiDet
Normal pedestal and noise with different HV for outer sensor of FST-18.
Normal pedestal and noise with different HV for inner sensor of FST-31.

02/19/2021:FST-28 Outer Sensor and FST-33 Inner Sensor RO test at SiDet
Inspect FST-28 and FST-33.
For FST 28, the wire-bonding between GND and Bias-GND close to APV 8 has some materials and is not bonded to the pad of Si Sensor shown below:

For FST-33, the VDD and GND is very close for APV 4 shown below:

Normal pedestal and noise with different HV for outer sensor of FST-28.
Normal pedestal and noise with different HV for inner sensor of FST-33.

02/22/2021
:FST-34 Outer Sensor, FST-35 Inner Sensor and FST-37 APV RO test at SiDet
Inspect FST-34, FST-35 and FST-37.
For FST 34, No visible broke wire. No visible floating wire. There is some transparent gel material on the  pad of sensor under the wire bonding of GND and bias-GND close to APV 5 shown below.

For FST 35, No visible broke wire. No visible floating wire.
For FST 37, the issue about the wire 13 and 14 is bonded the same pad is fixed. No visible broke wire. No visible floating wire. The distance between wire 14 and 15 is closer than others shown below.

Normal pedestal and noise with different HV for outer sensor of FST-34.
Normal pedestal and noise with different HV for inner sensor of FST-35.
Normal pedestal and noise with different HV for APV of FST-37.

02/23/2021
:FST-28 Inner Sensor and FST-33 Outer Sensor RO test at SiDet
Inspect FST-28 and FST-33.
For FST-28, No visible broke wire. No visible floating wire.
For FST-33, No visible broke wire. No visible floating wire. The VDD and GND is very close for APV 4 shown below (Fermilab technicians think it is fine).

For FST-28, There are an issue for the RO test of inner and out sensor when HV at 0 V. There are normal pedestal and noise when test inner and outer separately when HV at 0.
Normal pedestal and noise with different HV for outer sensor of FST-33.

02/24/2021:FST-28 Inner Sensor,  FST-34 Inner Sensor and FST-35 Outer Sensor RO test at SiDet
Inspect FST-34 and FST-35.
For FST 34, No visible broke wire. No visible floating wire.
For FST 35, No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for outer sensor of FST-35, and inner sensors of FST-34 and FST-28(For HV at 0, Normal pedestal and noise after ramping 140V/70V to 0V).
02/25/2021:FST-38 Inner Sensor RO test at SiDet
Inspect FST-38.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with HV 0 V and 70 V for inner sensor of FST-38. The current is too high for inner sensor when HV at 140 V(under the investigation).
02/25/2021 Open & Short Test Investigation
Test FST-51 
& FST-54.
No open and short connection found.
Place FST-51 in Storage Box #13 & FST-54 in Storage Box #14.
Place FST-51 & FST-54 in the transfer area in SiDet.

02/26/2021:FST-24 Inner Sensor,  FST-30 Outer Sensor and FST-37 Outer Sensor RO test at SiDet
Inspect FST-24, FST-30 and FST-37.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for inner sensor of FST-24.
Normal pedestal and noise with different HV for outer sensor of FST-30.
Normal pedestal and noise with different HV for outer sensor of FST-37.

03/01/2021:FST-38 Inner Sensor,  FST-51 APV and FST-54 APV RO test at SiDet
Inspect FST-51and FST-54The cooper layer of GND on outer sector was removed by Fermilab Technician.
No short found. 
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for APV of FST-51.
Normal pedestal and noise with different HV for APV of FST-54.
Remade the FST-38 inner Sensor RO test.
Normal pedestal and noise with different HV for inner sensor of FST-38.

03/02/2021:FST-37 Inner Sensor and FST-24 Outer Sensor RO test at SiDet
Inspect FST-24 and FST-37.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for inner
 sensor of FST-37.
Normal pedestal and noise with different HV for outer sensor of FST-24.
03/02/2021 Open & Short Test Investigation
Test FST-53
.
No open and short connection found.
Place FST-41 in Storage Box #15 & FST-42 in Storage Box #16 & FST-45 in Storage Box #17 & FST-53 in Storage Box #18.
Place FST-41 & FST-42 in the transfer area in SiDet.

03/03/2021
:FST-51 Inner Sensor and FST-54 Outer Sensor RO test at SiDet
Inspect FST-51 and FST-54.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for inner
 sensor of FST-51.
Normal pedestal and noise with different HV for outer sensor of FST-54.
Place FST-51 & FST-54 in the transfer area in SiDet.
03/03/2021 Open & Short Test Investigation
Test FST-48 & FST-39 & FST-55 & FST-17 & FST-27 & FST-52 & FST-25
.
No open and short connection found.
Place FST-48 in Storage Box #19 & FST-47 in Storage Box #20 & FST-36 in Storage Box #21 & FST-50 in Storage Box #22.

03/04/2021:FST-41 and FST-42 APV RO test at SiDet
Inspect FST-41 and FST-42.
No visible broke wire. No visible floating wire.
No short between VSS & GND, the resistance is ~9kohm. No short between VDD & GND. No short between SUB & GND. No short between SUB & VSS.
Normal pedestal and noise with different HV for APV
 of FST-41.
Normal pedestal and noise with different HV for APV of FST-42.
Place FST-41 & FST-42 in the transfer area of SiDet for sensor mount.
03/04/2021:Check purple cable
Check FST-V3.1-06 using prototype module. There is the
 I2C error. 

03/05/2021:FST-51 Outer Sensor and FST-54 Inner Sensor RO test at SiDet
Inspect FST-51 and FST-54.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for outer
 sensor of FST-51.
Normal pedestal and noise with different HV for inner sensor of FST-54.
03/05/2021:Check purple cable
Check FST-V3.1-04 and  
FST-V3.1-08 using prototype module with different HV. Normal behavior.
Check FST-V3.1-06 using prototype module. There is also the I2C error. 

03/09/2021:FST-42 Outer Sensor,  FST-38 Inner Sensor and FST-41 Inner Sensor RO test at SiDet
Inspect FST-41 and FST-42.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for outer
 sensor of FST-42.
Normal pedestal and noise with different HV for inner sensor of FST-41.
Remade the FST-38 inner Sensor current test.
Normal current for inner sensor of FST-38 when HV at 140 V.

03/15/2021:FST-45 and FST-53 APV, FST-42 Inner Sensor and FST-41 Outer Sensor RO test at SiDet
Inspect FST-41FST-42, FST-45 and FST-53.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for outer
 sensor of FST-41.
Normal pedestal and noise with different HV for inner sensor of FST-42.
Normal pedestal and noise with different HV for APV of FST-45 and FST-53.

03/16/2021: FST-30 Inner Sensor and FST-38 Outer Sensor RO test at SiDet
Inspect FST-30 and FST-38.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for inner
 sensor of FST-30.
Normal pedestal and noise for outer sensor of FST-38 when HV at 0V. There is a current issue when HV at 35V.
Place FST-22 in Storage Box #23 & FST-39 in Storage Box #24 & FST-43 in Storage Box #25 & FST-55 in Storage Box #26.

03/17/2021: FST-47 and FST-48 APV, FST-45 Inner Sensor and FST-53 Outer Sensor RO test at SiDet
Inspect FST-47, FST-48, FST-45 and FST-53.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for outer sensor of FST-53.
Normal pedestal and noise with different HV for inner sensor of FST-45.
Normal pedestal and noise with different HV for APV of FST-47 and FST-48.

03/18/2021:FST-36 and FST-50 APV RO test at SiDet
Inspect FST-36 and FST-50.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for APV of FST-36 and FST-50.

03/19/2021:
 FST-48 and FST-53 Inner Sensor, and FST-45 and FST-47 Outer Sensor RO test at SiDet
Inspect FST-47, FST-48, FST-45 and FST-53.
No visible broke wire. No visible floating wire. There is a scratch above C26 between APV 7 and APV 8 on outer sensor of FST-47 
Normal pedestal and noise with different HV for outer sensor of FST-45.
Normal pedestal and noise with different HV for inner sensor of FST-48 and FST-53.
Normal pedestal and noise with HV 70 V for outer sensor of FST-47.
There is a current issue when HV at 140 V 
for outer sensor of FST-47.

03/22/2021:
 FST-22 and FST-39 APV, FST-36 Inner Sensor and FST-50 Outer Sensor RO test at SiDet
Inspect FST-22, FST-39, FST-36 and FST-50.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for outer sensor of FST-50.
Normal pedestal and noise with different HV for inner sensor of FST-35.
Normal pedestal and noise with different HV for APV of FST-22 and FST-39.

03/23/2021:
 FST-17, FST-27, FST-43 and FST-55 APV RO test at SiDet
Inspect FST-17, FST-27, FST-43 and FST-55.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for APV of FST-17, FST-27, FST-43 and FST-55.

03/23/2021 Open & Short Test Investigation
Test FST-56 
& FST-58.
No open and short connection found.
Place FST-56 in Storage Box #29 & FST-58 in Storage Box #30.

03/24/2021: FST-22 and FST-50 Inner Sensor, FST-39 and FST-36 Outer Sensor RO test at SiDet
Inspect FST-22, FST-50, FST-39 and FST-36.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for outer sensor of FST-39 and FST-36.
Normal pedestal and noise with different HV for inner sensor of FST-22 and FST-50.

03/25/2021: FST-17 and FST-43 Inner Sensor, FST-27, FST-47, FST-38 and FST-55 Outer Sensor and FST-51 Sensor RO test at SiDet
Inspect FST-17, FST-43, FST-27 and FST-55.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for outer sensor of FST-27 and FST-55.
Normal pedestal and noise with different HV for inner sensor of FST-17 and FST-43.
Remade FST-51 sensor RO test due to encapsulation. Normal pedestal and noise with different HV.
Remade FST-47 outer sensor RO test. 
Normal current, pedestal and noise when HV at 140V.
Remade FST-38 outer sensor RO test. There is still a current issue when ramping HV from 0V to 35V.
03/26/2021: FST-56 and FST-58 APV RO, FST-39 Inner Sensor and FST-22 Outer Sensor, FST-51 RO test at SiDet
Inspect FST-56, FST-58, FST-39 and FST-22.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for outer sensor of FST-22.
Normal pedestal and noise with different HV for inner sensor of FST-39.
Normal pedestal and noise with different HV for APV of FST-56 and FST-58.
Remade FST-51 sensor RO test due to finishing inner sector encapsulation and repairing the broken wires of outer sector. 
Normal pedestal and noise with different HV.

03/29/2021: FST-51, FST-28,  FST-33, FST-24, FST-34, FST-35 and FST-30 Sensor RO test at SiDet
Remade FST-51 sensor RO test due to finishing outer sector encapsulation.  Normal pedestal and noise with different HV.
Remade 
sensor RO test for FST-35, FST-28, FST-33, FST-24, FST-34 and FST-30 due to starting the encapsulation.
Normal pedestal and noise with different HV for
sensor of FST-28, FST-33, FST-24, FST-34 and FST-30.
There is a current issue when HV ramping from 70 V to 105 V for FST-35 outer sensor.
Inspect FST-35, FST-28, FST-33, FST-24, FST-34 and FST-30.
No visible broke wire. No visible floating wire.
There is a long tail for oner wire of VDD on APV1 for FST-28.
Gave FST-22, FST-39, FST-33, FST-24, FST-34 and FST-30 to Fermilab technician for encapsulation.

03/30/2021: FST-27 and FST-55 Inner Sensor, FST-43 Outer Sensor, FST-28 and FST-37 RO test at SiDet
Inspect FST-27, FST-55, FST-43 and FST-17.
No visible broke wire. No visible floating wire.
There is one outer bond missing for APV 5 and APV 8 of FST-17.
Normal pedestal and noise with different HV for outer sensor of FST-43.
Normal pedestal and noise with different HV for inner sensor of FST-27 and FST-55.
Remade 
sensor RO test for FST-37 due to starting the encapsulation. Normal pedestal and noise with different HV.
Remade sensor RO test for FST-28 due to taking off a long tail for one wire of VDD on APV1. Normal pedestal and noise with different HV.

03/31/2021 Open & Short Test Investigation
Test FST-60 
& FST-61 & FST-62 & FST-63 & FST-64 & FST-65 & FST-66 & FST-72.
No open and short connection found.
Place FST-60 ~ FST-66 in Storage Box #31 ~ #37 & FST-72 in Storage Box #38.

04/01/2021: FST-22, FST-30,  FST-33, FST-34, FST-39, FST-24 and FST-54 Sensor RO test at SiDet
Remade sensor RO test for FST-22, FST-30, FST-33, FST-34, FST-39, FST-24 after the encapsulation. Normal pedestal and noise with different HV.
Remade sensor RO test for FST-54 due to starting the encapsulation. Normal pedestal and noise with different HV.
Gave FST-28, FST-37, FST-43, FST-55, FST-27 and FST-54 to Fermilab technician for encapsulation.

04/05/2021: FST-41, FST-42,  FST-45, FST-53, FST-35 and FST-36 Sensor RO test, and FST-17 and FST-47 Outer Sensor RO test at SiDet
Remade sensor RO test for FST-41, FST-42, FST-45 and FST-53 due to starting the encapsulation. Normal pedestal and noise with different HV. The current is higher than that of other modules for FST-45 inner sensor.
Remade sensor RO test for FST-35 and FST-36. There is still a current issue when HV ramping from 70 V to 105 V for FST-35 outer sensor. There is a current issue when HV ramping from 35 V to 70 V for FST-36 outer sensor.
Remade outer sensor RO test for FST-47. Normal pedestal and noise with different HV.
Made outer sensor RO test for FST-17. Normal pedestal and noise with different HV.

04/06/2021 Open & Short Test Investigation
Test FST-67 
& FST-68 & FST-69 & FST-70 & FST-71 & FST-75 & FST-76 & FST-77 & FST-78 & FST-79.
No open and short connection found.
There is a short connection between VDD and pin34 for FST-75 inner sector.
Place FST-67 ~ FST-71 in Storage Box #39 ~ #43 & FST-52 in Storage Box #44 & FST-76 ~ FST-79 in Storage Box #45 ~ #48.

04/07/2021: FST-28, FST-27,  FST-55, FST-43, FST-54, FST-37 and FST-48 Sensor RO test at SiDet
Remade sensor RO test for FST-28, FST-27, FST-55, FST-43, FST-54, FST-37 after the encapsulation. Normal pedestal and noise with different HV.
Inspect FST-48 and FST-47.
No visible broke wire and No visible floating wire for FST-48.
The wires of VDD are un-straight on APV 1 for FST-47.
There is one outer bond missing for APV 5 and APV 8 of FST-17.
Made outer 
sensor RO test for FST-48. Normal pedestal and noise when HV at 70V and 140V. These is a noise issue when HV at 0V.

04/08/2021: FST-17, FST-41, FST-42, FST-53, FST-47 and FST-50 Sensor RO test at SiDet
Remade sensor RO test for FST-17, FST-41, FST-42 and FST-53 after the encapsulation. Normal pedestal and noise with different HV.
Made sensor RO test for FST-47 after straightening the un-straight wires of VDD on APV1. Normal pedestal and noise when HV at 70 V and 140 V. There is a noise issue when HV at 0 V.
Remade sensor RO test for FST-50. There is still a noise issue when HV at 0 V.

04/09/2021: FST-61, FST-62 and FST-63 APV, and FST-56 inner Sensor RO test, and FST-38 current test at SiDet
Inspect FST-61, FST-62, FST-63, FST-56 and FST-58.
No visible broke wire. No visible floating wire.
The cooling tube of FST-58 is shorter than other modules. 
Normal pedestal and noise with different HV for inner sensor of FST-56.
Normal pedestal and noise with different HV for APV of FST-61, FST-62 and FST-63.
Tested the current 
with different HV after removing the two HV bias wires on the outer sensor relating to APV 5 and APV 6 for FST-38. There is a current issue when ramping HV from 105V to 140V for the first test. There is a current issue when ramping HV from 0V to 35V for the second test.

04/12/2021: FST-64, FST-65, FST-66 and FST-75 APV RO test at SiDet
Inspect FST-64, FST-65, FST-66 and FST-75.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for APV of FST-64, FST-65, FST-66 and FST-75.

04/15/2021: FST-61 and FST-62 inner, FST-63 and FST-56 outer, and FST-47, FST-48 and FST-50 Sensor RO test, FST-38 current test at SiDet
Inspect FST-61, FST-62, FST-63 and FST-56.
No visible broke wire. No visible floating wire.
Normal pedestal and noise with different HV for inner sensor of FST-61 and FST-62.
Normal pedestal and noise with different HV for outer sensor of FST-63 and FST-56.
Remade sensor RO test for FST-47, FST-48 and FST-50 after the encapsulation. Normal pedestal and noise when HV at 70V and 140V.
Tested the current with different HV after removing the two HV bias wires on the outer sensor relating to APV 7 and APV 8 for FST-38. There is still a current issue when ramping HV from 0V to 35V.